OpenStudy (anonymous):

1/7^(x-1)=1/2401 What is x?

OpenStudy (ranga):

2401 = 7^4 can you solve now?

OpenStudy (ranga):

\[\large \frac{ 1 }{ 7^{x-1} } = \frac{ 1 }{ 2401 } = \frac{ 1 }{ 7^4 }\]

OpenStudy (anonymous):

I ended up with x=-3 but my teacher told me my answer needed to be positive.

OpenStudy (ranga):

\[\large \frac{ 1 }{ 7^{x-1} } = \frac{ 1 }{ 7^4 }\] The only way that is possible is when x - 1 = 4 (because both bases are 7. The only way they can be equal is when the exponents are the same.) solve x - 1 = 4

OpenStudy (anonymous):

ok! i see where I went wrong! Thank you!

OpenStudy (ranga):

you are welcome.